### v1

# thanks: Dedicated to Prof. Jafar Sadeghi in honor of his 60th jubilee

Abstract
In this paper, we study the universe in the context of $f(T)$ gravity with flat-FRW background and consider it to be a combination of three components: baryonic matter, dark matter, and dark energy. We consider the new agegraphic as an alternative for dark energy and the Bose-Einstein condensation (BEC) as an alternative for dark matter. After that, first we obtain the Friedman equations and then we obtain the continuity equations in the presence of the interaction term between the dark components of the universe, so that the interaction term is considered as the flow of energy from dark matter to dark energy. In what follows, we plot the variation of the cosmological parameters of dark energy in terms of the redshift parameter by using of the power-law cosmology. Finally, we investigate the evolution and stability of the universe and report the values of the density parameters of the universe components that confirm the present model with observational data.
Equation of state parameter; Dark energy; Dark matter; Bose-Einstein condensate; $f(T)$ gravity.
I Introduction
Around the beginning of the new century, Riess et al [1] found out that the present universe is in an accelerated phase using type-Ia supernovae data, and then other groups [2, 3, 4] reconfirmed the same conclusion in the context of the cosmic microwave background radiation and the large-scale structure of the universe. The origin of these studies led to the definition of a mysterious energy called dark energy, which is responsible for the expansion of the universe. For this purpose, many models were studied in the fundamental theoretical framework such as string theory and quantum gravity, some of which can be mentioned such as the cosmological constant, modified gravity and teleparallel gravity [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47].
One of these models is the teleparallel gravity model, which was first proposed by Einstein to unify electromagnetism and gravity. In this theory, space-time is characterized by a linear connection without curvature and with twist by a metric tensor field called the dynamical tetrad field. Therefore, the geometry of general relativity, which is based on the Levi-Civita connection for a torsion-free curvature connection, gives way to the geometry of the parallel trap theory, which is based on the linear Weitzenböck connection in a curvature-free torsion connection. [48, 49, 50, 51, 53, 52, 54, 55].
On the other hand, another dark part of the universe is dark matter, which is not visible in any spectrum of electromagnetic radiation, it is an important topic in cosmology. Gravitational lensing and the cosmic microwave background are observational evidence of dark matter, which are calculated from the difference between the mass calculated for the giant celestial bodies [56, 57, 58, 59]. Therefore, this experimental evidence shows that dark matter causes the gravitational effect between the formation and evolution of galaxies and clusters and large-scale structures. For this purpose, we use the Bose-Einstein condensation model to understand the concept of dark matter. Bose-Einstein condensation occurs in boson-diluted gases, where a large number of bosons simultaneously occupy a ground quantum state in a system. A number of papers have addressed the connection between BEC and dark matter. Therefore, we follow the BEC model as one of the models that can investigate dark matter [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 77, 79, 80, 81, 82].
Therefore, in this work, we want to examine the evolution of the universe in the context of $f(T)$ gravity with equivalence BEC and the new agegraphic, respectively, with dark matter and dark energy. This study became the motivation to reach a new approach to understanding the universe by reconstructing this model in terms of the redshift parameter and replacing the dark parts with BEC and the new agegraphic.
This article is set up as follows:
In Sec. II, we imply the fundamental formation of $f(T)$ gravity in the flat-FRW metric. In Sec. III, we consider an interacting model between the dark parts of the universe. In Sec. IV, we review the model of the new agegraphic. In Sec. V, we consider BEC as dark matter and obtain a equation of state (EoS) for dark matter. In Sec. VI, we reconstruct the model in terms of the redshift parameter, and correspond the dark energy part with the new agegraphic. Finally, in Sec. VII, we will present a brief explanation for the present model.
II $f(T)$ gravity background
Here, we intend to review modified teleparallel gravity by $f(T)$ gravity background. Therefore, instead of the Levi-Civita connection in general relativity, we consider the Weitzenböck connection in teleparallel gravity. For this purpose, the corresponding geometry for each point $x^{\mu}$ in a manifold is defined based on the tetrad field $e_{i}(x^{\mu})$ as the basis vectors at same point, which will have $e_{i}.e_{j}=\eta_{ij}$, where the sign of Minkowski spacetime is $\eta_{ij}=diag(+1,-1,-1,-1)$. However, we can convert metric tensor $g_{\mu\nu}$ to the tetrad field $e^{i}_{\,\,\mu}$ as $g_{\mu\nu}=\eta_{\mu\nu}e^{i}_{\,\,\mu}e^{i}_{\,\,\nu}$, which the relations $e_{i}\,^{\mu}e^{i}\,_{\nu}=\delta_{\nu}^{\mu}$ and $e_{i}\,^{\mu}e^{j}\,_{\mu}=\delta_{i}^{j}$ are used. Note that Greek and Latin letters represent space-time components and tangent space-time components, respectively.
Nevertheless, the Weitzenböck connection is written down as
$\Gamma^{\lambda}_{~{}\mu\nu}=e_{i}^{~{}\lambda}\partial_{\mu}e^{i}_{~{}\nu}=-e^{i}_{~{}\mu}\partial_{\nu}e_{i}^{~{}\lambda},$ (1)
based on this, the torsion tensor ${T^{\lambda}}_{\mu\nu}$, the torsion scalar $T$, the antisymmetric tensor $S_{\lambda}^{\mu\nu}$, and the contortion tensor $K_{\lambda}^{\mu\nu}$ are
${T^{\lambda}}_{\mu\nu}=\Gamma^{\lambda}_{~{}\mu\nu}-\Gamma^{\lambda}_{~{}\nu\mu}={e_{a}}^{\lambda}\,\left({{\partial_{\mu}}{e^{a}}_{\nu}-{\partial_{\nu}}{e^{a}}_{\mu}}\right),$ (2a)
$T={S_{\lambda}}^{\mu\nu}\,{T^{\lambda}}_{\mu\nu},$ (2b)
$S_{\lambda}^{\mu\nu}=\frac{1}{2}\,\left({K_{\lambda}^{\mu\nu}+\delta_{\lambda}^{\mu}\,T_{\alpha}^{\alpha\nu}-\delta_{\lambda}^{\nu}\,T_{\alpha}^{\alpha\mu}}\right),$ (2c)
$K_{\lambda}^{\mu\nu}=-\frac{1}{2}\,\left({{T^{\mu\nu}}_{\lambda}-T_{\lambda}^{\nu\mu}-T_{\lambda}^{\mu\nu}}\right),$ (2d)
Now in this paper, we intend to explore $f(T)$ gravity model by the following action
$S=\int e\left(\frac{f(T)}{2\kappa^{2}}+\mathcal{L}_{m}\right)d^{4}x,$ (3)
where are $\kappa^{2}=8\pi G$ and $e=det\left(e^{i}_{\,\,\mu}\right)$, also $\mathcal{L}_{m}$ is the matter Lagrangian density. As we know in physics, the principle of least action is the variation principle of a system, so that by varying the above action with respect to a the tetrad field, we can obtain the field equation as
$e^{-1}\partial_{\mu}\left(e\,e^{\lambda}_{a}\,S_{\lambda}^{~{}\mu\nu}\right)f_{T}-e_{a}^{\rho}\,T^{\lambda}_{~{}\mu\rho}S_{\lambda}{}^{\nu\mu}f_{T}+e^{\lambda}_{a}S_{\lambda}{}^{\mu\nu}\,\partial_{\mu}(T)f_{TT}+\frac{1}{4}e_{a}^{\nu}f=\frac{1}{2}\kappa^{2}e_{a}^{\lambda}\mathcal{T}_{\lambda}{}^{\nu},$ (4)
where the indices $T$ and $TT$ indicate the first and second order derivatives with respect to torsion scalar, and $\mathcal{T}_{\lambda}{}^{\nu}$ is the energy-momentum tensor. Therefore, the Ricci tensor $R_{\mu\nu}$ and the Ricci scalar $R$ obtain in the following form
$R_{\mu\nu}=-\nabla^{\rho}S_{\nu\rho\mu}-g_{\mu\nu}\nabla^{\rho}T^{\sigma}_{~{}\rho\sigma}-S^{\rho\sigma}_{~{}~{}\mu}K_{\sigma\rho\nu},$ (5a)
$R=-T-2\nabla^{\mu}\left(T^{\nu}_{~{}\mu\nu}\right).$ (5b)
In next section, we explore an interacting model in context of $f(T)$ gravity in Friedmann-Robertson-Walker (FRW) metric.
III Interacting model in $f(T)$ gravity background
According to the previous section, we continue the corresponding system with regard to the homogeneity and isotropy of the universe. For this purpose, we take the background manifold as flat-FRW background as
$ds^{2}=dt^{2}-a^{2}(t)\left(dx^{2}+dy^{2}+dz^{2}\right),$ (6)
where $a(t)$ is the scale factor. The tetrad field is obtained as $e_{\mu}^{a}=\mathrm{diag}(1,a,a,a)$ or $e^{\mu}_{a}=\mathrm{diag}(1,a^{-1},a^{-1},a^{-1})$, and then the torsion scalar becomes:
$T=-6H^{2},$ (7)
where $H=\dot{a}/{a}$ is the Hubble parameter. The energy-momentum tensor is obtained as follows:
$\mathcal{T}_{i}^{j}=(\rho_{tot}+p_{tot})u_{i}u^{j}-p_{tot}\,\delta_{i}^{j},$ (8)
where $\rho_{tot}$ and $p_{tot}$ are respectively the total energy density and the total pressure of fluid inside the universe. Since 4-velocity is a 4-vector, one is parallel to the direction of the time, i.e., $u^{i}$ = (+1,0,0,0), so we will have $u_{i}u^{j}$ = 1. In that case, the non-zero elements of energy-momentum tensor are obtained in the following form
$\mathcal{T}_{i}^{j}=diag\left(\rho_{tot},-p_{tot},-p_{tot},-p_{tot}\right).$ (9)
Now according to the above, we can obtain the Friedmann equations as
$\rho_{tot}=\frac{6}{\kappa^{2}}H^{2}\partial_{T}f+\frac{1}{2\kappa^{2}}f,$ (10a)
${p}_{tot}=-\frac{2}{\kappa^{2}}\left(\dot{H}+3H^{2}\right)\partial_{T}f+\frac{24}{\kappa^{2}}H^{2}\dot{H}\partial_{TT}f-\frac{1}{2\kappa^{2}}f,$ (10b)
where the dot represents the derivative with respect to time evolution. The interesting point is that for the special case of $f(T)=T$, the above Friedman equation is the same as the standard Friedman equation in general relativity.
In what follows, we consider the content of the universe to be dominated by dark matter, baryonic matter and dark energy, the energy density and pressure of the universe can be written as three components of dark matter, baryonic matter and dark energy as follows:
$\rho_{tot}=\rho_{bm}+\rho_{dm}+\rho_{de},$ (11a)
${p}_{tot}=p_{bm}+p_{dm}+p_{de},$ (11b)
where indices $bm$, $dm$ and $de$ are introduced as the contents of baryonic matter, dark matter and dark energy. Hence, we can obtain the total continuity equation from Eqs. (10) as
$\dot{\rho}_{tot}+3H\left(\rho_{tot}+{p}_{tot}\right)=0,$ (12)
where continuity equations for the universe components with the presence of an interaction term between only dark energy and dark matter are obtained as follows:
$\dot{\rho}_{bm}+3H\left(\rho_{bm}+p_{bm}\right)=0,$ (13a)
$\dot{\rho}_{dm}+3H\left(\rho_{dm}+p_{dm}\right)=Q,$ (13b)
$\dot{\rho}_{de}+3H\left(\rho_{de}+{p}_{de}\right)=-Q,$ (13c)
where $Q$ is the interaction term. The interaction term appears in the universe when there is a flow of energy between the components of the universe, so that the direction of the energy flow is from dark energy to dark matter and there is no interaction between baryonic matter and other components. With this approach, the exclusive interaction between dark matter and dark energy without the involvement of visible matter leads to fascinating cosmological consequences. Note that the interaction term is confirmed from the point of view of the second law of thermodynamics and its value is positive [83]. From the continuity equation, we find that the unit of term $Q$ is equal to the product of the Hubble parameter in the energy density. Therefore, in this work, we take the corresponding interaction term as $Q=3b^{2}H\rho_{dm}$ in which $b$ is intensity of the energy flow transfer between the corresponding components of the universe. From Eqs. (13a) and (13b) we find
${\rho}_{bm}=\rho_{bm_{0}}a^{-3(1+\omega_{bm})},$ (14)
where $\omega_{bm}=p_{bm}/\rho_{bm}$ is Equation of State (EoS) of baryonic matter which is a constant, and $\rho_{bm_{0}}$ is the present dark energy of baryonic matter.
After that, we obtain dark energy contribution components from Eqs. (10) and (11) as
$\kappa^{2}\rho_{de}=6H^{2}\partial_{T}f+\frac{1}{2}f-\kappa^{2}\rho_{bm}-\kappa^{2}\rho_{dm},$ (15a)
$\kappa^{2}{p}_{de}=-2\left(\dot{H}+3H^{2}\right)\partial_{T}f+24H^{2}\dot{H}\partial_{TT}f-\frac{1}{2}f-\kappa^{2}\omega_{bm}\rho_{bm}-\kappa^{2}\omega_{dm}\rho_{dm},$ (15b)
where $\omega_{m}=p_{m}/\rho_{m}$ is as the matter EoS parameter which its value is a constant. Also, we can write down the EoS of dark energy as follows:
$\omega_{de}=\frac{p_{de}}{\rho_{de}}=-1-\frac{2\dot{H}\partial_{T}f-24H^{2}\dot{H}\partial_{TT}f+\kappa^{2}\rho_{m}\,(1+\omega_{m})}{6H^{2}\partial_{T}f+\frac{1}{2}f-\kappa^{2}\rho_{m}},$ (16)
where the obtained EoS of dark energy is a modified form, this means that if we choose $f(T)=T$ without matter, it becomes $\omega=-1-\frac{2\dot{H}}{3H^{2}}$, which is same as the standard form in general relativity. Then, the obtained EoS shows us that the nature of the universe depends on $f(T)$ gravity and matter.
IV The new agegraphic dark energy model
In this section, we give a general overview of the subject of the agegraphic. As we know the basis of quantum mechanics is based on the fact that things that can be measured have a special meaning in physics. Therefore, experimental observations can be the result of thoughts related to the specific science. In that case, the measurement of physical quantities in general relativity and quantum mechanics is different from each other in the sense that in the former, quantities are measured without any restrictions, but in the latter, the measurement is done with restrictions and discretely. Nevertheless, the theory of gravity that incorporates the principles of quantum mechanics is known as the quantum theory of gravity. Although the theory of quantum gravity is not fully established, some concepts based on the principles of quantum gravity are presented. This means that when gravitational fields such as the inflation period and the nature of black holes and quantum effects in gravity are very strong, they can be investigated using quantum gravity theory. Therefore, the phenomenology of space-time quantum fluctuations, which is explored by the theory of quantum gravity, is introduced as an agegraphic model. Accordingly, using the concept of space-time quantum fluctuations, the distance $t$ (with the speed of light $c=1$) in Minkowski space-time was determined according to the uncertainty principle with the following accuracy
$Eq16$ (17)
δ t=λ t_p^2/3t^1/3, where $t_{{p}}$ is Planck’s reduced time and $\lambda$ is a dimensionless constant of order unity [84, 85, 86, 87, 88]. The energy density obtained from the agegraphic model is equal to $\rho_{{DE}}\approx{\frac{1}{{\kappa}^{2}{t}^{2}}}$, where $t$ is the age of the universe. By substituting the age of the universe $t$ in the energy density of the agegraphic dark energy model with conformal time $\tau$, the energy density of the new agegraphic dark energy model is obtained as follows:
$\tau=\int{\frac{{dt}}{a}=\int_{0}^{a}{\frac{{da}}{H{a^{2}}}}},$ (18)
where $\dot{\tau}=\frac{1}{a}$. Therefore, the energy density of the new agegraphic dark energy reads
${\rho_{\Lambda}}=\,\frac{{3{n^{2}}}}{{{\kappa^{2}}{\tau^{2}}}},$ (19)
where $n^{2}$ is constant. In the next section, we will explore BEC model in the context of cosmological evolution.
V BEC model as dark matter
In this section, we intend to study the nature of dark matter as one of the dark universe parts by the regime of BEC. BEC is a state of matter in which the dilute Bose gas is cooled to such a low temperature that a phase transition occurs at this point and the macroscopic quantum phenomenon occurs as a quantum minimum state. Since dark matter has a significant contribution to the evolution of the universe, it leads us to adopt the BEC approach instead of dark matter. This means that dark matter is a type of BEC boson gas that dominates a part of the universe. Therefore,we consider the number density of particles by the Bose-Einstein statistics, which these particles are formed by decoupling of the remaining plasma in the early universe. However, the energy density of dark matter is equal to the mass of dark matter multiplied by the number density of particles, and dark matter pressure is defined by Bose-Einstein statistics in a sphere with a radius of particles momentum (refer to [89, 90, 91, 92, 93] for more details). In that case, by the generalized Gross-Pitaevskii equation [94, 95], we write down the dark matter pressure in terms of the energy density of dark matter as
$p_{dm}=\omega_{dm}\rho_{dm}^{2},$ (20)
where $\omega_{dm}=2\pi l_{a}/m_{dm}^{3}$, in which $l_{a}$ is the s-wave scattering length and here is introduced as EoS of dark matter. By inserting Eq. (20) into Eq. (13b), we have
$\rho_{dm}=\frac{1-b^{2}}{(1-b^{2})a^{3(1-b^{2})}-\omega_{dm}},$ (21)
where by inserting $a=a_{0}=1$ ($a_{0}$ is the current scale factor) we obtain the current $\rho_{dm_{0}}$ as
$\rho_{dm_{0}}=\frac{1-b^{2}}{1-b^{2}-\omega_{dm}},$ (22)
in that case, $\rho_{dm_{0}}$ reads
$\rho_{dm}=\rho_{dm_{0}}\,\frac{1-b^{2}-\omega_{dm}}{(1-b^{2})a^{3(1-b^{2})}-\omega_{dm}},$ (23)
where this relationship shows that the energy density of dark matter depends on the scale factor, interacting term, and BEC dark matter.
VI Reconstructing and correspondence of the model
In this section, we study the present model based on reconstructing and correspondence methods. In that case, we reconstruct the Friedmann equation in terms of redshift parameter, and then we correspond the agegraphic dark energy with $f(T)$ gravity. For this purpose, we write down the transformation between the scale factor and redshift parameter $z$ as $a(t)=a_{0}/(1+z)$ in which $a_{0}=1$ is the present scale factor, also differential form is as
$\frac{d}{dt}=-H(1+z)\frac{d}{dz}.$ (24)
Now in order to solve the model we consider a specific model called the power-law universe in the following form
$a(t)={a_{0}}{\left(\frac{t}{t_{0}}\right)^{m}},$ (25)
where $t_{0}$ is the present age of the universe, and $m$ is dimensionless positive coefficient. The Hubble’s parameter and consequently the age of the universe will be:
$H=\frac{m}{t},$ (26a)
$t_{0}=\frac{m}{H_{0}},$ (26b)
where $H(t_{0})=H_{0}=67.4\pm 0.5\,km\,s^{-1}\,Mpc^{-1}$ [96]. In that case, we can obtain the Hubble parameter in terms of $z$ as
$H(z)=H_{0}(1+z)^{\frac{1}{m}}.$ (27)
The interesting point is that we can calculate the value $m$ by fitting Eq. (27) with the astronomical data. For this purpose, we adopt the value of the coefficient $m$ from the result of Ref. [44] as $m=0.956$, from which we can immediately obtain the present age of the universe from Eq. (26b) as $t_{0}=13.75Gyr$. In what follows, we consider that the source of dark energy arises from the $f(T)$ gravity and the new agegraphic. Therefore, we correspond Eq. (19) with energy density of dark energy $\rho_{de}$ and obtains as
$\rho_{DE}\equiv\rho_{de}=\frac{3{n^{2}}(1-m)^{2}H_{0}^{2}}{\kappa^{2}\,m^{2}}\,(1+z)^{\frac{2(1-m)}{m}},$ (28)
where is used from relations
$T=\frac{6m^{2}}{t^{2}},$ (29a)
$\dot{H}=-\frac{m}{t^{2}},$ (29b)
$\tau=\int_{0}^{t}{\frac{{dt}}{a}}=\frac{m^{m}}{(1-m)H_{0}^{m}}\,t^{1-m}.$ (29c)
Now, by entering equations (14), (23), (27), and (28)in Eq. (15a), we can obtain the form of the function $f(T(z))$ including the value of $b=0.807$ as follows:
$f(z)=53.068n^{2}(1+z)^{-1.138}-\frac{1.395\kappa^{2}\rho_{bm_{0}}}{2\omega_{bm}+1.303}(1+z)^{3\omega_{bm}+1.954}$
$+\kappa^{2}\rho_{dm_{0}}\left(\omega_{dm}-0.349\right)\left(11.737\ln\!\left(1+z\right)-5.735\ln\!\left(-0.349(1+z)+\omega_{dm}(1+z)^{2.046}\right)\right),$ (30)
where by using Eqs. (7) and (27), which results in $z=-1+\left(-T/H_{0}^{2}\right)^{n/2}$, we convert $f(z)$ as the form of the function $f(T)$ (it is unnecessary to write it here). In what follows, we can obtain the dark energy pressure $p_{de}$ by inserting Eq. (28) into the continuity equation (13c), and then we obtain the varition of $\rho_{de}$, $p_{de}$, and $\omega_{de}$ for dark energy component in terms of redshift parameter as shown in Figs. 1 and 2. We note that the free parameters has the key role for dynamical description of the cosmological parameters. In that case, we choice them based on positivity the dark energy density and negativity the dark energy pressure for the late universe ($z=0$). Therefore, the free parameters are selected as $n=18$, $\omega_{dm}=2.5$, $\rho_{dm_{0}}=4873$, $\omega_{bm}=1$, and $\rho_{bm_{0}}=545$ which $n$, $\omega_{dm}$, and $\omega_{bm}$ are dimensionless quantities, and $\rho_{dm_{0}}$ and $\rho_{bm_{0}}$ are dimension $M^{4}$ in Planck units. We can see in Figs. 1 that the value of energy density decrease from a high positive value in the early universe to a lower positive value in the late universe ($z=0$), also the value of pressure decrease from a high positive value in the early universe to a lower negative value in the late universe ($z=0$).
Figure 1: the variations of $\rho_{de}$ and $p_{de}$ in terms of the redshift parameter.
We note that the EoS parameter is a very important parameter so that can understand at a glance how the dynamics of the universe are evolving from the early time to the late time. Therefore, we notice that the matter period and the accelerated pase period for the universe can be expressed as $0<\omega<\frac{1}{3}$ and $-1<\omega<-\frac{1}{3}$, respectively. Hence, Fig. 2 shows us that the universe begins from matter era to accelerated phase era in which the value of EoS is about $-0.71$ in the present time ($z=0$), that represent the universe is undergoing an accelerated expansion.
Figure 2: the variations of $\omega_{de}$ in terms of the redshift parameter.
In what follows, we explore the stability of the model by using the sound speed parameter which one comes from the concept of thermodynamics. For this reason, we consider the universe to be an adiabatic thermodynamic system where no heat or mass is transferred out, meaning that the perturbation has zero entropy. Therefore, the sound speed parameter $c_{s}^{2}$ is defined in the following form
$c_{s}^{2}=\frac{\partial p_{de}}{\partial\rho_{de}}=\frac{\partial_{z}p_{de}}{\partial_{z}\rho_{de}},$ (31)
where the index $z$ indicates derivative with respect to redshift parameter. We now draw $c_{s}^{2}$ in terms of the redshift parameter as shown in Fig. 3. We note that the conditions of larger and smaller sound speed parameters, i.e., $c_{s}^{2}>0$ and $c_{s}^{2}<0$, indicate stability and instability, respectively. In that case, Fig. 3 displays that the universe begins from a stability phase and then at the threshold of the accelerated phase enters a non-stability phase in late time. Therefore, this variation represents that the energy density of dark energy in the late universe is not in a controlled growth.
Figure 3: The graph of the sound speed (dimensional quantity) in terms of redshift parameter by $\alpha=50$, $\beta=0.001$, $b=0.01$, $c=2$, $m=0.95$, $r=0.5$, $g=-1$, $\lambda=0.5$, $a_{0}=0.1$, and $\xi=0.25,1.25,2.25$.
At the end, we test the present model by using the density parameters of dark energy, dark matter, and baryonic matter. For this purpose, we account the corresponding density parameters as follows:
$\Omega_{bm_{0}}=\frac{\rho_{bm_{0}}}{\rho_{c}},$ (32a)
$\Omega_{dm_{0}}=\frac{\rho_{dm_{0}}}{\rho_{c}},$ (32b)
$\Omega_{de_{0}}=\frac{\rho_{de_{0}}}{\rho_{c}},$ (32c)
where $\rho_{c}=3H^{2}/\kappa^{2}$ is the critical density, and index zero is introduced the present values of the corresponding quantity. Now, by substituting the relationships related to energy density (14), (23) and (28) in the above relationship, the contribution of the universe components in the late universe is calculated according to table 1. The contents of Table 1 are calculated for the late time state ($z=0$), and show that the matter-energy contribution into the universe is consistent with the observational data [96]. The energy budget values from this reference are derived from precise measurements of the anisotropies of the cosmic microwave background (CMB) and other cosmological observations.
Table 1: The present values of the density parameters of $\Omega_{bm_{0}}$, $\Omega_{dm_{0}}$, and $\Omega_{de_{0}}$.
 $\rm{the~{}density~{}parameter}$ $\rm{the~{}present~{}value}$
$\Omega_{bm_{0}}$ $0.04$
$\Omega_{dm_{0}}$ $0.27$
$\Omega_{de_{0}}$ $0.69$
VII Conclusion
In this paper, we explored the universe in the context of $f(T)$ gravity by flat-FRW background. We considered the universe to be a combination of three components called baryonic matter, dark matter and dark energy. For this purpose, we took the origin of dark matter from BEC and considered the source of dark energy from the new agegraphic. After calculating Einstein’s equation and then Friedman’s equations for $f(T)$ gravity, we wrote down the continuity equations separately for the above three components. Also, we considered an interaction between the dark components of the universe, such that energy flows from dark matter to dark energy. In what follows, we considered the BEC to be a bosonic dilute gas material that, if cooled to a very low temperature, reaches the quantum ground state. Therefore, we obtained the EoS of BEC as a dark matter as $p_{dm}=\omega_{dm}\rho_{dm}^{2}$. On the other hand, we corresponded the energy density of $f(T)$ gravity with the energy density of new agegraphic. Next, we took the power-law cosmology for the scale factor, which the Hubble parameter has been obtained in terms of the redshift parameter. Then, we plotted the variation of $\rho_{de}$, $p_{de}$, and $\omega_{de}$ in terms of the redshift parameter which showed that the universe is undergoing accelerated expansion. Also, we explored stability analysis of the current model by the sound speed parameter, which the acquired results represented the universe is in a non-stability phase at the late universe. Finally, we test our model by calculating the present values of the density parameters of $\Omega_{bm_{0}}$, $\Omega_{dm_{0}}$, and $\Omega_{de_{0}}$, which indicate the universe is consistent with the observational data.